Evaluate the line integral: βˆ«π’–(π‘₯,𝑦,𝑧)×ⅆ𝒓,


where 𝒖(π‘₯,𝑦,𝑧)=(𝑦^2,π‘₯,𝑧) and the curve π‘ͺ is described by 𝒛=𝑦=𝑒^π‘₯ with π‘₯∈[0,1].


1
Expert's answer
2021-11-28T16:52:51-0500

"\int {u(x,y,z)dr = } \int\limits_C {{y^2}} dx + xdy + zdz"

"y = {e^x} \Rightarrow dy = {e^x}dx"

"z = {e^x} \Rightarrow dz = {e^x}dx"

Then

"\int\limits_C {{y^2}} dx + xdy + zdz = \int\limits_0^1 {{e^{2x}}dx + x{e^x}} dx + {e^x} \cdot {e^x}dx = \int\limits_0^1 {2{e^{2x}}dx + } \int\limits_0^1 {x{e^x}} dx = \left. {{e^{2x}}} \right|_0^1 + \int\limits_0^1 {x{e^x}} dx = {e^2} - {e^0} + \int\limits_0^1 {x{e^x}} dx"

Since

"\int\limits_0^1 {x{e^x}} dx = \left| {\begin{matrix}
{u = x}&{dv = {e^x}dx}\\
{du = dx}&{v = {e^x}}
\end{matrix}} \right| = \left. {x{e^x}} \right|_0^1 - \int\limits_0^1 {{e^x}dx = } \left. {x{e^x}} \right|_0^1 - \left. {{e^x}} \right|_0^1 = e - 0 - e + {e^0} = 1"

Then

"\int\limits_C {{y^2}} dx + xdy + zdz = {e^2} - {e^0} + \int\limits_0^1 {x{e^x}} dx = {e^2} - 1 + 1 = {e^2}"